OpenStudy (anonymous):

y = –2x² + 2 y = –x² y = 2x² y = 3x² – 4

OpenStudy (hoblos):

so...

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

graph please

OpenStudy (anonymous):

Place values in x and y so that when x = 0, y = ? or when y = 0, x = ? So that you can plot.

OpenStudy (anonymous):

for the first one

OpenStudy (mysesshou):

http://www.wolframalpha.com/input/?i=plot+y+%3D+%E2%80%932x%5E2+%2B+2+ (If the link doesn't work..) If you type "plot y = –2x^2 + 2 " into the question space, it will plot it for you to check your answers. +2 is the Y intercept. Then use Order's method of plug and chug for values of x and y.

OpenStudy (anonymous):

y=-2x^2 when y = 0 0=-2x^2+2 0=-22x^2 x^2=1 x=1 (1,0) When x=0 y=-2(0)^2+2 y=2 (0,2) Those are the points on the graph of the first equation

OpenStudy (anonymous):

 for the first equation